### v1

# Leveraging Nested MLMC for Sequential Neural Posterior Estimation with Intractable Likelihoods

Abstract
Sequential neural posterior estimation (SNPE) techniques have been recently proposed for dealing with simulation-based models with intractable likelihoods. They are devoted to learning the posterior from adaptively proposed simulations using neural network-based conditional density estimators. As a SNPE technique, the automatic posterior transformation (APT) method proposed by Greenberg et al. (2019) performs notably and scales to high dimensional data. However, the APT method bears the computation of an expectation of the logarithm of an intractable normalizing constant, i.e., a nested expectation. Although atomic APT was proposed to solve this by discretizing the normalizing constant, it remains challenging to analyze the convergence of learning. In this paper, we propose a nested APT method to estimate the involved nested expectation instead. This facilitates establishing the convergence analysis. Since the nested estimators for the loss function and its gradient are biased, we make use of unbiased multi-level Monte Carlo (MLMC) estimators for debiasing. To further reduce the excessive variance of the unbiased estimators, this paper also develops some truncated MLMC estimators by taking account of the trade-off between the bias and the average cost. Numerical experiments for approximating complex posteriors with multimodal in moderate dimensions are provided.
1 Introduction
Simulator-based models are widely used across various scientific disciplines, including neuroscience [47], physics [7, 24], biology [10, 29, 39, 46], and inverse graphics [52]. These models serve as crucial tools for describing and comprehending investigated processes based on observed data. However, when applying traditional Bayesian inference to simulator-based models, challenges arise, such as the intractable likelihood function $p(x|\theta)$ and the computational expense associated with running the simulator.
To address these challenges, a series of likelihood-free Bayesian computation (LFBC) methods have been developed. These methods include approximate Bayesian computation (ABC) [5, 45], synthetic likelihoods (SL) [50, 57], Bayes optimization [27], likelihood-free inference by ratio estimation [55], and pseudo marginals methods [2, 3]. A comprehensive summary and review of these methods can be found in [11], and they have all been benchmarked in [32, 41].
Posterior density estimation approaches approximate the posterior of interest $p(\theta|x_{o})$ with a family of density estimators $q_{\phi}(\theta)$, where $\phi$ is the parameters of the density estimators. Optimization-based approaches are widely used in these methods, and the Kullback-Leibler (KL) divergence between $p(\theta|x)$ and $q_{\phi}(\theta)$, which measures the differences between two densities, is commonly chosen as the loss function. Variational Bayes (VB), as a computationally optimization-based effective method for approximating the posterior distribution of a Bayesian problem, is widely used. In the likelihood-free context, Tran et al., [56] developed a new VB method with an intractable likelihood, while He et al., [30] proposed an unbiased VB method based on nested MLMC. However, these method tend to fail in cases where simulations are expensive, as the nested estimation in these methods requires additional simulation procedures.
Out of this reason, there has been a growing interest in employing neural networks to represent probability density recently, particularly normalizing flows [38]. When the inference problem is focused solely on the observation $x_{o}$, the data efficiency can be improved using the sequential training schemes proposed in the sequential neural posterior estimation (SNPE) methods [25, 42, 48]. In this approach, model parameters are drawn from a proposal distribution that is more informative about $x_{o}$ compared to the prior distribution. However, SNPE requires a modification of the loss function compared to neural posterior estimation (NPE) in order to ensure that the neural networks approximates the true posterior $p(\theta|x)$. Many approaches have been developed to address this issue.
Among the SNPE methods, automatic posterior transformation (APT) requires the computation of the expectation of the logarithm of an intractable normalizing constant, which is a nested expectation. They then propose to use the atomic APT to discretize the normalizing constant. However, it remains challenging to utilize the existing analysis techniques [6] to perform convergence analysis to the best of our knowledge. To address this limitation, we propose the nested APT method, an method enable the convergence analysis without scarifying the performance. We conduct a thorough analysis of the bias and variance of the loss estimators, moreover, we provide a convergence result for this biased estimator in the case of stochastic gradient descent (SGD). To evaluate the effectiveness of the nested APT, we conduct a series of comprehensive numerical experiments.
Since nested APT is a biased estimator, we undertake a comprehensive exploration of a range of unbiased methods based on multi-level Monte Carlo (MLMC) [17] to eliminate the bias. We establish several theorems that investigate the order of bias, variance, and the average cost of the estimators of the loss function and its gradient. Additionally, as we use stochastic gradient-based optimization methods to update the parameters of the neural network, it is well known that the unbiased method suffers from excessive variance in the gradient. To address this issue, we employ several variants of truncated MLMC methods to diminish the variance. Furthermore, we compute the order of the variance and average cost for the estimators for the gradient of the loss function. Finally, we utilize these results to provide the convergence analysis of truncated MLMC and nested APT.
The remainder of this paper is organized as follows. In Section 2, we propose the nested APT method and provide insights into possible failures when directly using the nested APT. Moreover, we propose a strategy to address this failure. We also present the experimental results and analyze them in this section. In Section 3, we introduce the basics of MLMC methods, including their formulation and theoretical analysis of the random sequence utilized in MLMC methods. Equipped with these tools, we develop both unbiased MLMC methods and truncated MLMC methods and analyze the order of variance for the losses and gradients, as well as the average cost. In Section 4, we provide convergence analysis for nested APT and truncated MLMC methods in the case of SGD. In Section 5, we conduct a series of numerical experiments on benchmark tasks. The results of these experiments are presented and analyzed in this section. Finally, we conclude this paper with some remarks in Section 6.
2 Nested APT
2.1 Problem formulation
Let $p(\theta)$ represent the prior distribution of the model parameter of interest. Given an observed sample $x_{o}$, our objective is to perform inference on the posterior distribution $p(\theta|x_{o})\propto p(\theta)p(x_{o}|\theta)$. However, in many cases, the likelihood function $p(x|\theta)$ either lacks an explicit expression or presents challenges in direct evaluation. Instead, it can be expressed using a ‘simulator’ approach, wherein, given a fixed model parameter $\theta$, we can sample $x$ from $p(x|\theta)$.
Since our objective is to approximate the posterior of interest $p(\theta|x_{o})$ using tractable density estimators, the KL divergence serves as a primary measure of the discrepancy between two densities. The KL divergence is defined as
$\mathcal{D}_{\mathrm{KL}}\left(p(\theta)\|q(\theta)\right)=\int p(\theta)\log\frac{p(\theta)}{q(\theta)}\mathrm{d}\theta,$
which is not less than 00 by using Jensen’s inequality and attains its minima when $q(\theta)$ agrees with $p(\theta)$, making it suitable as a loss function. When the likelihood $p(x|\theta)$ is tractable, one can directly approximate $p(\theta|x_{o})$ by minimizing the KL divergence between the target distribution and the proposed estimator $q(\theta)$ within a certain family of distributions. In the likelihood-free context, this can be viewed as a problem of conditional density estimation. Within this framework, a conditional density estimator $q_{F(x,\phi)}(\theta)$ based on a neural network [48, 25], is utilized to approximate $p(\theta|x)$ over the admissible set of tuning parameter $\phi\in\Phi$. To this end, we focus on minimizing the following average KL divergence under the marginal distribution $p(x)=\int p(\theta)p(x|\theta)\mathrm{d}\theta$
$\quad\ \mathbb{E}_{p(x)}\left[\mathcal{D}_{\mathrm{KL}}\left(p(\theta|x)\|q_{F(x,\phi)}(\theta)\right)\right]$
$=\iint p(x)p(\theta|x)\left(\log p(\theta|x)-\log q_{F(x,\phi)}(\theta)\right)\mathrm{d}x\mathrm{d}\theta$
$=-\mathbb{E}_{p(\theta,x)}\left[\log q_{F(x,\phi)}(\theta)\right]+\iint p(\theta,x)\log p(\theta|x)\mathrm{d}x\mathrm{d}\theta$
$:=\mathcal{L}(\phi)+\iint p(\theta,x)\log p(\theta|x)\mathrm{d}x\mathrm{d}\theta,$
where the term
$\mathcal{L}(\phi):=-\mathbb{E}_{p(\theta,x)}\left[\log q_{F(x,\phi)}(\theta)\right],$ (1)
is used as the loss function. However, since $\mathcal{L}(\phi)$ is intractable, we use its empirical estimator instead
^$\hat{\mathcal{L}}_{N}(\phi)=-\frac{1}{N}\sum_{i=1}^{N}\log q_{F(x_{i},\phi)}(\theta_{i}),$ (2)
where the training data $\{(\theta_{i},x_{i})\}_{i=1}^{N}$ is sampled from the joint probability density $p(\theta,x)=p(\theta)p(x|\theta)$. After training, given the observation $x_{o}$, the posterior $p(\theta|x_{o})$ can be approximated by $q_{F(x_{o},\phi)}(\theta)$.
Since our aim is to conduct conditional density estimation at $x_{o}$, it is essential to utilize a proposal distribution $\tilde{p}(\theta)$ that provides more informative priors regarding $x_{o}$ in comparison to the prior distribution $p(\theta)$. After initializing $\tilde{p}(\theta)$ as $p(\theta)$, we then want the approximation of $p(\theta|x_{o})$ to serve as a good proposal in the following simulation. This conditional density estimation with adaptively chosen proposal is called sequential neural posterior estimation (SNPE). However, after replacing $p(\theta,x)$ with $\tilde{p}(\theta,x)=\tilde{p}(\theta)p(x|\theta)$ in (1), it is observed that $q_{F(x,\phi)}(\theta)$ approximates the proposal posterior:
$\tilde{p}(\theta|x)=p(\theta|x)\frac{\tilde{p}(\theta)p(x)}{p(\theta)\tilde{p}(x)},$ (3)
where $\tilde{p}(x)=\int\tilde{p}(\theta)p(x|\theta)\mathrm{d}\theta$. Hence, we need to adjust loss function $\mathcal{L}(\phi)$ to make $q_{F(x,\phi)}(\theta)$ approximate the true posterior $p(\theta|x)$.
In APT [25], the proposal distribution $\tilde{p}(\theta)$ is initialized as the prior distribution $p(\theta)$. Consequently, (1) can be directly used for the loss function. In the subsequent rounds, [25] proposed to replace $q_{F(x,\phi)}(\theta),\ p(\theta,x)$ with $\tilde{q}_{F(x,\phi)}(\theta)$, $\tilde{p}(\theta,x)$ in (1) respectively. Explicitly, the loss function proposed in APT is
$\mathbb{E}_{\tilde{p}(\theta,x)}\left[-\log\tilde{q}_{F(x,\phi)}(\theta)\right],$ (4)
where
$\tilde{q}_{F(x,\phi)}(\theta)$ $=q_{F(x,\phi)}(\theta)\frac{\tilde{p}(\theta)}{p(\theta)}\frac{1}{Z(x,\phi)},$ (5)
$Z(x,\phi)$ $=\int\frac{q_{F(x,\phi)}(\theta^{\prime})}{p(\theta^{\prime})}\tilde{p}(\theta^{\prime})\mathrm{d}\theta^{\prime},$
$Z(x,\phi)$ here denotes the normalizing constant. Proposition 1 in [48] shows that if $q_{F(x,\phi)}(\theta)$ is expressive enough that $\tilde{q}_{F(x,\phi^{*})}(\theta)=\tilde{q}(\theta)$ for some parameter $\phi^{*}$, then $q_{F(x,\phi^{*})}(\theta)=q(\theta|x)$.
Since the integral $Z(x,\phi)$ is usually intractable in practice, APT proposes the use of ‘atomic’ proposals, APT with such proposal is known as atomic APT. Specifically, they assume that $\tilde{p}(\theta)=U_{\Theta}$, where $U_{\Theta}$ is a uniform distribution over a finite set $\Theta=\{\theta_{1},\dots,\theta_{M}\}$. The uniform setting of the proposal distribution allows for the analytical computation of $Z(x,\phi)$. With this method, (3) and (5) can be reformulated as
$\tilde{p}(\theta|x)=\frac{p(\theta|x)/p(\theta)}{\sum_{\theta^{\prime}\in\Theta}p(\theta^{\prime}|x)/p(\theta^{\prime})},\quad\tilde{q}_{x,\phi}(\theta)=\frac{q_{F(x,\phi)}(\theta)/p(\theta)}{\sum_{\theta^{\prime}\in\Theta}q_{F(x,\phi)}(\theta^{\prime})/p(\theta^{\prime})}.$ (6)
Proposition 1 in [25] provides the consistency guarantees of atomic APT: given that each $\Theta$ is constructed by sampling $\theta^{\prime}$ from a distribution that covers the target $p(\theta|x_{o})$, atomic APT is able to recover the full posterior.
However, to the best of our knowledge, the use of atomic proposals makes it challenging to analyze its convergence behaviour. Therefore unable to explain the unexpected low performance in some tasks [13] with the existing convergence results [6], both in terms of insight and theory. As an alternative approach to estimate $Z(x,\phi)$, the nested estimation, which enjoys a comprehensive theoretical framework and comparable performance, is studied in the next section.
2.2 Nested APT method
For ease of presentation, we denote $g_{\phi}(x,\theta):=q_{F(x,\phi)}(\theta)/p(\theta)$. We reformulate (4) as
$\quad\ \mathbb{E}_{\tilde{p}(\theta,x)}\left[-\log\tilde{q}_{F(x,\phi)}(\theta)\right]$
$=-\mathbb{E}_{\tilde{p}(\theta,x)}\left[\log q_{F(x,\phi)}(\theta)+\log\frac{\tilde{p}(\theta)}{p(\theta)}\right]+\mathbb{E}_{\tilde{p}(x)}[\log Z(x,\phi)]$
$=-\mathbb{E}_{\tilde{p}(\theta,x)}\left[\log\frac{q_{F(x,\phi)}(\theta)}{p(\theta)}\right]+\mathbb{E}_{\tilde{p}(x)}[\log Z(x,\phi)]-\mathbb{E}_{\tilde{p}(\theta)}\left[\log\tilde{p}(\theta)\right]$
$=-\mathbb{E}_{\tilde{p}(\theta,x)}\left[\log g_{\phi}(x,\theta)\right]+\mathbb{E}_{\tilde{p}(x)}\left[\log\mathbb{E}_{\tilde{p}(\theta^{\prime})}\left[g_{\phi}(x,\theta^{\prime})\right]\right]-\mathbb{E}_{\tilde{p}(\theta)}\left[\log\tilde{p}(\theta)\right]$
:=~$:=\tilde{\mathcal{L}}(\phi)-\mathbb{E}_{\tilde{p}(\theta)}\left[\log\tilde{p}(\theta)\right],$
where
~$\tilde{\mathcal{L}}(\phi)=-\mathbb{E}_{\tilde{p}(\theta,x)}\left[\log g_{\phi}(x,\theta)\right]+\mathbb{E}_{\tilde{p}(x)}\left[\log\mathbb{E}_{\tilde{p}(\theta^{\prime})}\left[g_{\phi}(x,\theta^{\prime})\right]\right],$ (7)
is selected as the loss function, $-\mathbb{E}_{\tilde{p}(\theta)}\left[\log\tilde{p}(\theta)\right]$ in the last column is dropped for it is independent of $\phi$. We employ stochastic gradient methods to optimize the loss function in this paper, therefore $\nabla_{\phi}\tilde{\mathcal{L}}(\phi)$ is of interest. By allowing the interchange of expectation and gradient operator, the gradient of (7) is given by
$\nabla_{\phi}\tilde{\mathcal{L}}(\phi)=-\mathbb{E}_{\tilde{p}(\theta,x)}\left[\nabla_{\phi}\log g_{\phi}(x,\theta)\right]+\mathbb{E}_{\tilde{p}(x)}\left[\nabla_{\phi}\log\mathbb{E}_{\tilde{p}(\theta^{\prime})}\left[g_{\phi}(x,\theta^{\prime})\right]\right].$ (8)
Notice that (8) holds a similar formulation of the gradient use in the optimization procedure of Bayesian experimental design [9, 16, 34, 37]. Given any query point $(\theta,x)\sim\tilde{p}(\theta,x)$, we denote the corresponding query of the loss and gradient as
$\psi_{\phi}$ $=-\log g_{\phi}(x,\theta)+\log Z(x,\phi),$ (9)
$\rho_{\phi}$ $=-\nabla_{\phi}\log g_{\phi}(x,\theta)+\nabla_{\phi}\log Z(x,\phi),$ (10)
so that $\mathbb{E}_{\tilde{p}(\theta,x)}[\psi_{\phi}]=\tilde{\mathcal{L}}(\phi)$ and $\mathbb{E}_{\tilde{p}(\theta,x)}[\rho_{\phi}]=\nabla_{\phi}\tilde{\mathcal{L}}(\phi)$. With the abuse of notation aiming at easing representation, unless otherwise specified in the following paper, $\mathcal{L}(\phi)$ and $\nabla_{\phi}\mathcal{L}(\phi)$ will be used to represent (7) and (8) respectively.
Given the intractability of the normalizing constant $Z(x,\phi)=\mathbb{E}_{\tilde{p}(\theta^{\prime})}\left[g_{\phi}(x,\theta^{\prime})\right]$, it remains an obstacle to derive an estimator for the loss function. A simple choice is leveraging its empirical estimator based on $M$ samples, which is given by
$\hat{Z}_{M}(x,\phi)=\frac{1}{M}\sum_{j=1}^{M}g_{\phi}(x,\theta^{\prime}_{j}),$ (11)
where $\theta^{\prime}_{1},\cdots,\theta^{\prime}_{M}\sim\tilde{p}(\theta^{\prime})$ independently, $M$ in the subscript denotes the number of the samples used for the estimation of $Z(x,\phi)$. We thus arrive at queries for nested estimators of (9) and (10) respectively
$\psi_{\phi,M}$ $=-\log g_{\phi}(x,\theta)+\log\hat{Z}_{M}(x,\phi)=\log\frac{1}{M}\sum_{j=1}^{M}\frac{g_{\phi}(x,\theta^{\prime}_{j})}{g_{\phi}(x,\theta)},$ (12)
$\rho_{\phi,M}$ $=-\nabla_{\phi}\log g_{\phi}(x,\theta)+\nabla_{\phi}\log\hat{Z}_{M}(x,\phi).$ (13)
Nested estimators for the loss function (7) and its gradient (8) are then given by the mean of $N$ iid copies of their queries
^$\hat{{\mathcal{L}}}^{\mathrm{Ne}}(\phi)$ $=\frac{1}{N}\sum_{i=1}^{N}\psi_{\phi,M}^{(i)}=\frac{1}{N}\sum_{i=1}^{N}\log\frac{1}{M}\sum_{j=1}^{M}\frac{g_{\phi}(x_{i},\theta^{\prime}_{ij})}{g_{\phi}(x_{i},\theta_{i})},$ (14)
$\nabla_{\phi}\hat{\mathcal{L}}^{\text{Ne}}(\phi)$ $=\frac{1}{N}\sum_{i=1}^{N}\rho^{(i)}_{\phi,M},$ (15)
where $\theta_{ij}^{\prime}\stackrel{{\scriptstyle iid}}{{\sim}}\tilde{p}(\theta^{\prime})$ constitute the inner sample with a size of $M$ and $(\theta_{i},x_{i})\stackrel{{\scriptstyle iid}}{{\sim}}\tilde{p}(\theta,x)$ constitute the outer sample with a size of $N$, $\psi^{(i)}_{\phi,M},\ \rho^{(i)}_{\phi,M}$ are iid copies of $\psi_{\phi,M}$ and $\rho_{\phi,M}$ respectively. Due to the nonlinearity of the logarithm, (14) is a biased estimator. Using similar arguments in [53], it is not difficult to evaluate that the order of variance of the estimator of nested APT ^$\hat{{\mathcal{L}}}^{\mathrm{Ne}}(\phi)$ is of $O(1/N)$, and $O(1/M)$ for its bias.
2.3 Performance evaluations for comparing nested APT with atomic APT
In this section, we conduct a series of experiments to demonstrate the performance of nested APT compared to atomic APT. For detailed experimental setting, we refer the readers to Section 5. The results are reported in Figure 1, from which we observe that the proposed nested APT achieves comparable performance with atomic APT. A noteworthy advantage of our proposed nested APT method is that it enables the application of a series of existing result for an optimizer with a biased gradient.
Due to its biased nature, the nested APT method may not directly benefit from the existing variance reduction techniques, as the effects of variance and bias on the optimal gap are ‘tangled’. To address this issue, we propose to make use of the existing multi-level Monte Carlo (MLMC) techniques for the improvement of the nested estimator.
A
B
C
Figure 1: Density plot for nested APT and atomic APT. A. Two-moon model, from left to right: available ground truth, atomic APT with inner samples $M=100$, nested APT with inner samples $M=100$. B. Lotka-Volterra model, from left to right: ground truth simulated with SMC-ABC [4], atomic APT with $M=100$, nested APT with $M=100$. C. M/G/1 queue model, the setting are the same with Lotka-Volterra.
3 Unbiased multilevel nested APT
3.1 Basic idea of MLMC
Nested simulation combined with the MLMC method has been studied for other applications [18, 19, 22, 30]. In this paper, MLMC is used to derive an unbiased estimator for APT. Increasing $M$ for $\rho_{\phi,M}$ eventually recovers the desired $\rho_{\phi}$, this allows the exchange of expectation and limitation that
$\lim_{M\to\infty}\mathbb{E}\left[\rho_{\phi,M}\right]=\mathbb{E}\left[\rho_{\phi}\right]=\nabla_{\phi}\mathcal{L}(\phi).$ (16)
Then we take a monotonically increasing sequence for inner sample size $\{M_{\ell}\}_{\ell=1}^{\infty}$ with $M_{\ell}=2^{\ell}M_{0}=2M_{\ell-1}$, then (16) can be reformulated as
$\nabla_{\phi}\mathcal{L}(\phi)=\mathbb{E}\left[\rho_{\phi,M_{0}}\right]+\sum_{\ell=1}^{\infty}\mathbb{E}\left[\rho_{\phi,M_{\ell}}-\rho_{\phi,M_{\ell-1}}\right].$ (17)
Set $\{\Delta\rho_{\phi,\ell}\}_{\ell=0}^{\infty}$ to be a sequence of random variables satisfying
$\mathbb{E}\left[\Delta\rho_{\phi,0}\right]=\mathbb{E}\left[\rho_{\phi,M_{0}}\right],\quad\mathbb{E}\left[\Delta\rho_{\phi,\ell}\right]=\mathbb{E}\left[\rho_{\phi,M_{\ell}}-\rho_{\phi,M_{\ell-1}}\right]\quad(\ell\geq 1),$
then (17) can be reformulated in terms of the infinite sum of the expected values of $\{\Delta\rho_{\phi,\ell}\}_{\ell=0}^{\infty}$, i.e., $\nabla_{\phi}\mathcal{L}(\phi)=\sum_{\ell=0}^{\infty}\mathbb{E}\left[\Delta\rho_{\phi,\ell}\right]$. In the case of the loss function, we also introduce the similar sequence $\{\Delta\psi_{\phi,\ell}\}_{\ell=0}^{\infty}$ satisfying
$\mathbb{E}\left[\Delta\psi_{\phi,0}\right]=\mathbb{E}\left[\psi_{\phi,M_{0}}\right],\quad\mathbb{E}\left[\Delta\psi_{\phi,\ell}\right]=\mathbb{E}\left[\psi_{\phi,M_{\ell}}-\psi_{\phi,M_{\ell-1}}\right]\quad(\ell\geq 1),$ (18)
similarly we also have $\mathcal{L}(\phi)=\sum_{\ell=0}^{\infty}\mathbb{E}\left[\Delta\psi_{\phi,\ell}\right]$.
3.1.1 Antithetic construction
Motivated by the prevailing literature on MLMC [20, 8, 19, 22], this paper presents the construction of the sequence $\{\Delta\rho_{\phi,\ell}\}_{\ell=0}^{\infty}$ with antithetic coupling, leading to a faster rate of convergence for smooth functions. The fundamental concept of antithetic coupling involves the selection of two non-overlapping subsets, each with a size of $M_{\ell-1}$ from the $M_{\ell}$ inner samples $\{\theta_{j}^{\prime}\}_{j=1}^{M_{\ell}}$ utilized for computing $\rho_{\phi,M_{\ell}}$ given in (13). This results in two independent realizations of $\rho_{\phi,M_{\ell-1}}$, identified as $\rho^{(a)}_{\phi,M_{\ell-1}}$ and $\rho^{(b)}_{\phi,M_{\ell-1}}$. Specifically,
$\rho^{(a)}_{\phi,M_{\ell-1}}=-\frac{\nabla_{\phi}g_{\phi}(x,\theta)}{g_{\phi}(x,\theta)}+\frac{\nabla g^{(a)}_{\phi,M_{\ell-1}}(x)}{g^{(a)}_{\phi,M_{\ell-1}}(x)},$
where
$g^{(a)}_{\phi,M_{\ell-1}}(x)=\frac{1}{M_{\ell-1}}\sum_{j=1}^{M_{\ell-1}}g_{\phi}(x,\theta_{j}^{\prime}),\ \nabla g^{(a)}_{\phi,M_{\ell-1}}(x)=\frac{1}{M_{\ell-1}}\sum_{j=1}^{M_{\ell-1}}\nabla_{\phi}g_{\phi}(x,\theta_{j}^{\prime}).$
The notations $\rho^{(b)}_{\phi,M_{\ell-1}}$ , $\nabla g^{(b)}_{\phi,M_{\ell-1}}(x)$, and $g^{(b)}_{\phi,M_{\ell-1}}(x)$ are defined in the similar way by using $\theta^{\prime}_{j}$, $j=M_{\ell-1}+1,\dots,M_{\ell}$ instead. Then the antithetic construction of $\{\Delta\rho_{\phi,\ell}\}_{\ell=0}^{\infty}$ is then given as
$\Delta\rho_{\phi,0}=\rho_{\phi,M_{0}},\quad\Delta\rho_{\phi,\ell}=\rho_{\phi,M_{\ell}}-\frac{1}{2}\left(\rho^{(a)}_{\phi,M_{\ell-1}}+\rho^{(b)}_{\phi,M_{\ell-1}}\right)\quad(\ell\geq 1).$ (19)
The antithetic construction of $\{\Delta\psi_{\phi,\ell}\}_{\ell=0}^{\infty}$ can be similarly given.
3.1.2 Variance of multilevel Monte Carlo estimator
In this section, we will give a series of detailed evaluation of the variance of $\Delta\rho_{\phi,\ell},\rho_{\phi,M_{\ell}}$ and $\Delta\psi_{\phi,\ell}$ which will serve as important tools when we evaluate the variance of some other MLMC methods in the coming sections. Before we present the main results, we firstly introduce the following key lemma which is stated as Lemma 1 in [19].
Lemma 3.1.
Let $X$ be a real-valued random variable with mean zero, and let $\overline{X}_{N}$ be an average of $N$ i.i.d samples of $X$. If for $\mathbb{E}[|X|^{u}]<\infty$ for $u\geq 2$, there exists a constant $C_{u}>0$ depending only on $u$ such that
$\mathbb{E}\left[\left|\overline{X}_{N}\right|^{u}\right]\leq C_{u}\dfrac{\mathbb{E}[|X|^{u}]}{N^{u/2}},\quad\mathbb{P}\left[\left\lvert\overline{X}_{N}\right\rvert>c\right]\leq C_{u}\dfrac{\mathbb{E}\left[|X|^{u}\right]}{c^{u}N^{u/2}},$
for any $c>0$.
The following theorem studies the expectation and variance of $\Delta\psi_{\phi,\ell}$ as $\ell\to\infty$. In this paper, the notation $\lesssim$ is used to signifies the omission of a constant that is irrelevant to the asymptotic behavior.
Theorem 3.2.
If there exist $s,z>2$ with $(s-2)(z-2)\geq 4$ such that for any $\phi$,
$\quad\mathbb{E}_{x,\theta^{\prime}}\left[\left|\frac{g_{\phi}(x,\theta^{\prime})}{Z(x,\phi)}\right|^{s}\right]<\infty\quad and\quad\mathbb{E}_{x,\theta^{\prime}}\left[\left|\log\frac{g_{\phi}(x,\theta^{\prime})}{Z(x,\phi)}\right|^{z}\right]<\infty,$
we have
$\mathrm{E}\left[\|\Delta\psi_{\phi,\ell}\|_{2}^{2}\right]\lesssim M_{\ell}^{-r_{1}},$
where $M_{\ell}=M_{0}2^{\ell}$, $r_{1}=\min(s(z-2)/2z,2)$.
Proof.
This proof follows an argument similar to Theorem 2 in [22]. ∎
When stochastic gradient descent method is applied, it is also desirable to examine the expectation and variances of the $\ell_{2}$ norm of the gradient. To this end, we establish the following theorem based on the work of [23]. The proof of this theorem is detailed in Appendix A.1.
Theorem 3.3.
Assume that there exists $s>2$ such that for any $\phi$
$\underset{x,\phi,\theta}{\sup}\|\nabla_{\phi}\log g_{\phi}(x,\theta)\|_{\infty}<\infty,\,\mathbb{E}_{x,\theta}\left[\left|\frac{g_{\phi}(x,\theta)}{Z(x,\phi)}\right|^{s}\right]<\infty,$
we then have
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}\right]\lesssim M_{\ell}^{-1},\quad\mathrm{E}\left[\|\Delta\rho_{\phi,\ell}\|_{2}^{2}\right]\lesssim M_{\ell}^{-r_{2}},$
where $M_{\ell}=M_{0}2^{\ell}$, $r_{2}=\min(s,4)/2\in(1,2]$.
For the convergence of $\rho_{\phi,M_{\ell}}$, we establish the following theorem.
Theorem 3.4.
Under the setting of Theorem 3.3, we have $\mathrm{Var}\left[\rho_{\phi,M_{\ell}}\right]\lesssim M_{\ell}^{-1}.$
A detailed version and proof of this theorem is provided in Appendix A.2. This theorem can be similarly extended to the case where arbitrarily number of inner samples are used for the estimation of the gradient. As a result, it indicates the convergence rate of the gradient estimator is the same as the loss function.
3.2 Unbiased MLMC for APT
3.2.1 RU-MLMC
Recall that both loss function and its gradient can be represented as the summation of the expected value of the random variable sequence, and this topic has been well studied. We introduce a non-negative integer-valued random variable $L$ that independent of the two random variable sequences, with its probability mass function $\mathbb{P}(L=\ell)=w_{\ell}$. We then have
$\mathbb{E}\left[{\omega_{L}}^{-1}\Delta\rho_{\phi,L}\right]=\sum_{\ell=0}^{\infty}\mathbb{E}\left[{\omega_{\ell}}^{-1}\Delta\rho_{\phi,\ell}\right]P(L=\ell)=\nabla_{\phi}\mathcal{L}(\phi).$
The equivalence between $\nabla_{\phi}\mathcal{L}(\phi)$ and the expectation of the query
$V_{\mathrm{RU}}={\omega_{L}}^{-1}{\Delta\rho_{\phi,L}}$ (20)
leads to an unbiased Monte Carlo estimator of the gradient $\nabla_{\phi}\mathcal{L}(\phi)$. We similarly define $U_{\mathrm{RU}}={\omega_{L}}^{-1}{\Delta\psi_{\phi,L}}$ for the query of the estimation of the loss function $\mathcal{L}(\phi)$. This method is known as randomized unbiased multilevel Monte Carlo (RU-MLMC) method [51]. In this study, we take $L$ as a geometric distribution $\mathrm{Ge}(p)$ with $\omega_{\ell}=(1-p)^{\ell}p$ and $p=1-2^{-\alpha}$. To ensure a finite variance and finite expected computational cost for $V_{\mathrm{RU}},U_{\mathrm{RU}}$, it is required that $\alpha\in(1,\min\left(r_{1},r_{2}\right))$. This is further examined with the following theorem.
Theorem 3.5.
Under the setting of Theorems 3.2 and 3.3, let $L\sim\mathrm{Ge}(p)$ with $p=1-2^{-\alpha}$ and $1<\alpha<\min\left(r_{1},r_{2}\right)$, where $r_{1}$ and $r_{2}$ are from Theorems 3.2 and 3.3 respectively. Then we have
$\mathrm{Var}\left[U_{\mathrm{RU}}\right]$ $\leq\frac{A}{(1-2^{\alpha-r_{1}})(1-2^{-\alpha})},$
$\mathrm{Var}\left[V_{\mathrm{RU}}\right]$ $\leq\frac{B}{(1-2^{\alpha-r_{2}})(1-2^{-\alpha})},$
$\mathrm{Cost}_{\mathrm{RU}}$ $\propto M_{0}\frac{2^{\alpha}-1}{2^{\alpha}-2},$
where the constants $A,B$ are independent of $\alpha$ and $r_{1}$.
Proof.
The proof of this theorem follows in line of Theorem 3.6. ∎
This theorem indicates that a large $\alpha$ leads to less expected total computational burden, but larger variance for the estimators of loss and its gradient.
3.2.2 GRR-MLMC
The Russian roulette (RR) estimator [44] is also employed to estimate the sum of an infinite series, wherein the evaluation of any term in the series only demands a finite amount of computation. This estimator relies on randomized truncation and assigns a higher weight to each term to accommodate the possibility of not computing them. The query of the gradient of RR-MLMC estimator is given by
$V_{\mathrm{RR}}:=\sum_{j=0}^{L}\frac{\Delta\rho_{\phi,j}}{p_{j}},$ (21)
where $p_{j}=\mathbb{P}(L\geq j)=\sum_{\ell=j}^{\infty}w_{\ell}$ for $j\geq 0$. If all $p_{j}>0$, then $V_{\mathrm{RR}}$ is unbiased since
$\mathbb{E}[V_{\mathrm{RR}}]$ $=\sum_{\ell=0}^{\infty}\mathbb{P}(L=\ell)\sum_{j=0}^{\ell}\frac{\mathbb{E}[\Delta\rho_{\phi,j}]}{p_{j}}=\sum_{j=0}^{\infty}\frac{\mathbb{E}[\Delta\rho_{\phi,j}]}{p_{j}}\sum_{\ell\geq j}^{\infty}\mathbb{P}(L=\ell)$
$=\sum_{j=0}^{\infty}\mathbb{E}[\Delta\rho_{\phi,j}]=\nabla_{\phi}\mathcal{L}(\phi).$
Motivated by [43], in order to trade lower variance with higher cost based on the conventional RR-MLMC, one way is to ensure that the first ${\underline{m}}$ terms of the infinite series are always computed, where this ${\underline{m}}$ is called the base level. In this case, the random index $L$ is set to have a lower bound ${\underline{m}}$, i.e., $\mathbb{P}(L\geq{\underline{m}})=0$, implying $p_{j}=1$ for all $j\leq{\underline{m}}$. We call this modified estimator as generalized Russian roulette (GRR) estimator. When ${\underline{m}}=0$, this degenerates to the conventional RR-MLMC. The associated quaery for GRR-MLMC estimator of the gradient is
$V_{\mathrm{GRR}}:=\rho_{\phi,M_{{\underline{m}}}}+\sum_{j={\underline{m}}+1}^{L}\frac{\Delta\rho_{\phi,j}}{p_{j}}.$ (22)
In this case, the probability mass function of $L$ is chosen as $\mathbb{P}(L={\underline{m}})=1-\sum_{\ell>{\underline{m}}}w_{\ell}$ and $\mathbb{P}(L=\ell)=w_{\ell}$ for $\ell>{\underline{m}}$, where $w_{\ell}=(1-p)^{\ell}p$ and $p=1-2^{-\alpha}$. We denote this distribution by $\mathrm{Ge}(p,{\underline{m}})$. This implies that $V_{\mathrm{GRR}}$ is also unbiased since all $p_{j}>0$,
Similarly, we define the query for GRR estimator of the loss function:
$U_{\mathrm{GRR}}:=\psi_{\phi,M_{{\underline{m}}}}+\sum_{j={\underline{m}}+1}^{L}\frac{\Delta\psi_{\phi,j}}{p_{j}}.$ (23)
$U_{\mathrm{GRR}}$ and the following theorem is established aiming at evaluating their order.
Theorem 3.6.
Under the setting of Theorems 3.2 and 3.3, let $L\sim\mathrm{Ge}(p,{\underline{m}})$ with $p=1-2^{-\alpha}$ and $1<\alpha<\min\left(r_{1},r_{2}\right)$, where $r_{1}$ and $r_{2}$ are from Theorems 3.2 and 3.3, and the base level ${\underline{m}}\geq 1$, we have
$\mathrm{Var}\left[U_{\mathrm{GRR}}\right]$ $\leq A\left(\frac{1}{2^{{\underline{m}}}}+\frac{2^{(\alpha-r_{1})({\underline{m}}+1)}}{1-2^{\alpha-r_{1}}}\right),$
$\mathrm{Var}\left[V_{\mathrm{GRR}}\right]$ $\leq B\left(\frac{1}{2^{{\underline{m}}}}+\frac{2^{(\alpha-r_{2})({\underline{m}}+1)}}{1-2^{\alpha-r_{2}}}\right),$
$\mathrm{Cost}_{\mathrm{GRR}}$ $\propto M_{0}2^{{\underline{m}}}+M_{0}\frac{2^{(1-\alpha)({\underline{m}}+1)}}{1-2^{(1-\alpha)}},$
where the constants $A,B$ are independent of $\alpha,\ r_{1},\ r_{2},\ \underline{m}$.
The proof of this theorem is detailed in Appendix A.3.
3.3 Truncated MLMC for APT
When treating $\nabla_{\phi}\mathcal{L}(\phi)$ as the summation of the infinite random variable sequence $\{\Delta\rho_{\phi,\ell}\}_{\ell=0}^{\infty}$, we notice that $\Delta\rho_{\phi,0}=\rho_{\phi,0}$ is just the nested APT estimator with a bias of order $O(1/M_{0})$. For $\Delta\rho_{\phi,\ell}$ with $\ell\geq 1$, it actually contributes to the reduction of bias while simultaneously increasing the variance and this additional variance grows with $\ell$. An empirical demonstration of the additional variance for $\Delta\rho_{\phi,\ell}$ is shown in Figure 2.
Figure 2: Left: Crude RU-MLMC after the third round. Right: The truncated RU-MLMC ($\overline{m}=4$) after the third round.
We observe that the truncated RU-MLMC method enjoys a more stable loss compared to the original one, indicating that truncation helps reducing variance. Therefore, in order to mitigate the variance of the RU-MLMC, we propose to truncate the distribution of $L$ by setting the largest value of $L$ as $\overline{m}$. The truncation leads to a bias of order $O(1/M_{{\overline{m}}})$. Following [33], we extend the truncated idea to other unbiased MLMC methods.
3.3.1 TGRR-MLMC
TGRR-MLMC stands for the truncated version of GRR-MLMC. Let ${\overline{m}}\geq{\underline{m}}$ be a truncated level, and let $L$ be a nonnegative integer-value random index taking values in $\{{\underline{m}},\dots,{\overline{m}}\}$. The query for the estimation of the gradient of TGRR-MLMC is then defined as
$V_{\mathrm{TGRR}}=\rho_{\phi,M_{{\underline{m}}}}+\sum_{j={\underline{m}}+1}^{L}\frac{\Delta\rho_{\phi,j}}{p_{j}},$ (24)
where $p_{j}=\mathbb{P}(L\geq j)>0$ for ${\underline{m}}\leq j\leq{\overline{m}}$. This estimator is biased:
$\mathbb{E}[V_{\mathrm{TGRR}}]$ $=\mathbb{E}[\rho_{\phi,M_{{\underline{m}}}}]+\sum_{\ell={\underline{m}}+1}^{{\overline{m}}}\mathbb{P}(\tilde{L}=\ell)\sum_{j={\underline{m}}+1}^{\ell}\frac{\mathbb{E}[\Delta\rho_{\phi,j}]}{p_{j}}$
$=\mathbb{E}[\rho_{\phi,M_{{\underline{m}}}}]+\sum_{j={\underline{m}}+1}^{{\overline{m}}}\frac{\mathbb{E}[\Delta\rho_{\phi,j}]}{p_{j}}\sum_{\ell\geq j}^{{\overline{m}}}\mathbb{P}(\tilde{L}=\ell)$
$=\mathbb{E}[\rho_{\phi,M_{{\underline{m}}}}]+\sum_{j={\underline{m}}+1}^{{\overline{m}}}\mathbb{E}[\Delta\rho_{\phi,j}]=\mathbb{E}[\rho_{\phi,M_{{\overline{m}}}}].$
We should note that if ${\overline{m}}={\underline{m}}=m$, TGRR-MLMC degenerates to nested APT with the inner sample size $M_{m}=M_{0}2^{m}$. For any truncated method with truncated level $\overline{m}$, they all share the bias of order $\mathcal{O}(M_{\overline{m}})$ based on (45).
For TGRR-MLMC, the probability mass function of the random index $L$ is chosen as
$\mathbb{P}(L=\ell)=\begin{cases}1-\frac{\sum_{\ell={\underline{m}}+1}^{{\overline{m}}}w_{\ell}}{1-(1-p)^{{\overline{m}}+1}},\ &\ell={\underline{m}}\\ \frac{w_{\ell}}{1-(1-p)^{{\overline{m}}+1}},\ &{\underline{m}+1}\leq\ell\leq{\overline{m}}\\ 0,\ &\text{otherwise},\end{cases}$ (25)
where $w_{\ell}=(1-p)^{\ell}p$ and $p=1-2^{-\alpha}$. We denote this distribution by $\mathrm{Ge}(p,{\underline{m}},{\overline{m}})$. The whole procedure of TGRR-MLMC method is summarized in Algorithm 1.
Input: Prior $p(\theta)$, implicit simulator model $p(x|\theta)$, neural network $q_{F(x,\phi)}(\theta)$, $1<\alpha<\min(r_{1},r_{2})$, $M_{0},M_{\ell}=M_{0}2^{\ell}$, truncated level ${\overline{m}}$, base level ${\underline{m}}$, total round $R$, optimizer $g(\cdot,\cdot)$, batch size $B$, outer sample size $N$, $t=0$, dataset $\mathcal{D}_{\mathrm{out}}$ for outer samples, dataset $\mathcal{D}_{\mathrm{in}}(x)$ for the inner samples of $x$, $\tilde{p}_{1}(\theta)=p(\theta)$
1 for $k$ in $\{1,2,\cdots,R\}$ do Generate $\{(\theta_{i},x_{i})\}_{i=1}^{N}$ from $\tilde{p}_{k}(\theta)p(x|\theta)$ and level $\{\ell_{i}\}_{i=1}^{N}$ from $\mathrm{Ge}(p,{\underline{m}},{\overline{m}})$ Update dataset $\mathcal{D}_{\mathrm{out}}\leftarrow\{(\theta_{i},x_{i},\ell_{i})\}_{i=1}^{N}\cup\mathcal{D}_{\mathrm{out}}$
2 if $k=1$  then repeat Generate $\{(\theta_{i},x_{i})\}_{i=1}^{B}\sim\mathcal{D}_{\mathrm{out}}$ Update parameter $\phi_{t+1}\leftarrow g\left(\phi_{t},-\frac{1}{N}\sum_{i=1}^{N}\nabla_{\phi}\log q_{F(x_{i},\phi_{t})}(\theta_{i})\right)$ $t\leftarrow t+1$ until $\phi_{t}$ converged
3end if
4if $k\geq 2$ then repeat Generate $\{(\theta_{i^{\prime}},x_{i^{\prime}})\}_{i=1}^{B}\sim\mathcal{D}_{\mathrm{out}}$
5 for $i^{\prime}$ in $\{1,2,\cdots,B\}$ do Generate $\{\theta^{\prime}_{j}\}_{j=1}^{M_{{\underline{m}}}-1}$ from $\tilde{p}_{k}(\theta^{\prime})$ $\mathcal{D}_{\mathrm{in}}(x_{i^{\prime}})\leftarrow\{\theta^{\prime}_{j}\}_{j=1}^{M_{{\underline{m}}}-1}\cup\{\theta_{i^{\prime}}\}$ and compute $\rho^{(i^{\prime})}_{\phi_{t},M_{m}}$ with $\mathcal{D}_{\mathrm{in}}(x_{i^{\prime}})$ for $\ell^{\prime}$ in $\{\{{\underline{m}}+1,{\underline{m}}+2,\cdots,\ell_{i}\}$} do $\mathcal{D}^{(a)}_{\mathrm{in}}(x_{i^{\prime}})\leftarrow\{\theta^{\prime}_{j}\}_{j=1}^{M_{\ell^{\prime}-1}-1}\cup\{\theta_{i^{\prime}}\}$, $\mathcal{D}^{(b)}_{\mathrm{in}}(x_{i^{\prime}})\leftarrow\{\theta^{\prime}_{j}\}_{j=M_{\ell^{\prime}-1}}^{M_{\ell^{\prime}}-1}\cup\{\theta_{i^{\prime}}\}$ $\mathcal{D}_{\mathrm{in}}(x_{i^{\prime}})\leftarrow\mathcal{D}^{(a)}_{\mathrm{in}}(x_{i^{\prime}})\cup\mathcal{D}^{(b)}_{\mathrm{in}}(x_{i^{\prime}})$ Compute $\Delta\rho^{(i^{\prime})}_{\phi_{t},\ell^{\prime}}$ with $\mathcal{D}_{\mathrm{in}}(x_{i^{\prime}})$ using (19) end for
6Compute $V_{\mathrm{TGRR}}^{(i^{\prime})}$ using (24) end for
7Update parameter $\phi_{t+1}\leftarrow g\left(\phi_{t},-\frac{1}{B}\sum_{i^{\prime}=1}^{B}V_{\mathrm{TGRR}}^{(i^{\prime})}\right)$ $t\leftarrow t+1$ until $\phi_{t}$ converged
8end if
9Update proposal $\tilde{p}_{k+1}(\theta)\leftarrow q_{F(x_{o},\phi_{t})}(\theta)$ end for
Algorithm 1 Truncated Generalized Russian Roulett (TGRR)
4 Convergence results of SGD
In this section, we analyze the convergence of the methods proposed in this paper with some existing techniques. For simplicity, we only study the widely used SGD with constant step size $\gamma$ in this paper. Our result should extend to boarder cases. The parameter update procedure of SGD is presented as
$\phi_{t+1}=\phi_{t}-\gamma_{t}\nabla_{\phi}\hat{\mathcal{L}}(\phi_{t}),$
leveraging the information of an estimator of the gradient of the loss function at the current state $\phi_{t}$. The gradient estimator for TGRR-MLMC is
$\nabla_{\phi}\hat{\mathcal{L}}^{\text{TGRR}}(\phi):=\frac{1}{N}\sum_{i=1}^{N}V_{\mathrm{TGRR}}^{(i)},$
where $V_{\mathrm{TGRR}}^{(i)}$ are iid copies of $V_{\mathrm{TGRR}}$. The gradient estimators for other MLMC methods follows a similar form. The gradient estimator for nested APT is stated previously in (15). Since some of them are biased estimators for $\nabla_{\phi}\mathcal{L}(\phi)$, we take the following decomposition
$\nabla_{\phi}\hat{\mathcal{L}}(\phi)=\nabla_{\phi}\mathcal{L}(\phi)+b(\phi)+\eta(\phi),$ (26)
where $b(\phi)$ and $\eta(\phi)$ denote the bias and the noise of gradient estimator $\nabla_{\phi}\hat{\mathcal{L}}(\phi)$ respectively.
Assumption 4.1.
There exist constants $U_{b}<\infty$ and $U_{\eta}<\infty$ such that
$\underset{\phi\in\Phi}{\sup}\left\|b(\phi)\right\|_{2}^{2}$ $=\underset{\phi\in\Phi}{\sup}\left\|\mathbb{E}\left[\nabla_{\phi}\hat{\mathcal{L}}(\phi)\right]-\nabla_{\phi}\mathcal{L}(\phi)\right\|^{2}_{2}\leq U_{b},$
$\underset{\phi\in\Phi}{\sup}\ \mathbb{E}\left[\left\|\eta(\phi)\right\|_{2}^{2}\right]$ $=\underset{\phi\in\Phi}{\sup}\ \mathbb{E}\left[\left\|\nabla_{\phi}\hat{\mathcal{L}}(\phi)-\mathbb{E}\left[\nabla_{\phi}\hat{\mathcal{L}}(\phi)\right]\right\|_{2}^{2}\right]\leq U_{\eta}.$
Following the similar procedure to derive (45), we find that the order of $U_{b}$ for nested APT is $\mathcal{O}(1/M)$, while $\mathcal{O}(1/(M_{0})^{\overline{m}})$ for TGRR. Intuitively speaking, when it is set that $M_{0}^{\overline{m}}=M$ and $\underline{m}<\overline{m}$, $M$ inner samples are always required for each outer sample in nested APT, while the inner sample size for TGRR-MLMC is determined by the distribution $\mathrm{Ge}(p,\underline{m},\overline{m})$ whose maximal output is $M_{0}^{\overline{m}}=M$. Therefore, this truncated MLMC method takes less cost than nested APT to achieve the same level of bias.
We then focus on studying how $U_{b}$ and $U_{\eta}$ would affect the optimal gap at any $t>0$, which is defined as
$G_{t}=\mathbb{E}\left[\mathcal{L}(\phi_{t})\right]-\mathcal{L}(\phi^{*}),$ (27)
where $\phi^{*}=\underset{\phi\in\Phi}{\arg\min}\,\mathcal{L}(\phi)$. Moving forward, we introduce the following two commonly used assumptions when analyzing the convergence of SGD, as outlined in [6].
Assumption 4.2.
The objective function $\mathcal{L}(\phi)$ is differentiable and there exists a constant $K$ such that for every $\phi,\phi^{\prime}$
$\mathcal{L}(\phi)\leq\mathcal{L}(\phi^{\prime})+\nabla_{\phi}\mathcal{L}(\phi)^{T}(\phi-\phi^{\prime})+\frac{K}{2}\left\|\phi-\phi^{\prime}\right\|^{2}_{2}.$
Assumption 4.3.
The objective function $\mathcal{L}(\phi)$ is differentiable and there exists $\mu>0$ such that
$\left\|\nabla\mathcal{L}(\phi)\right\|_{2}^{2}\geq 2\mu(\mathcal{L}(\phi)-\mathcal{L}(\phi^{*})).$
In situations where a fixed number of iterations are exclusively employed for parameter updates, the evaluation of the impact of variance and bias on the upper bound of the optimal gap is undertaken through the application of the theorem presented below, following [1].
Theorem 4.4.
Under Assumptions 4.2 and 4.3 and the constant step size $\gamma\leq\min\left\{1/K,1/\mu\right\}$, for a fixed $T$ steps of SGD, the upper bound of the optimal gap (27) is given as
$G_{T}\leq(1-\gamma\mu)^{T}G_{0}+\frac{1}{2\mu}\left(U_{b}+U_{\eta}\right).$ (28)
Since $G_{0}$ depends solely on the initialization of $\phi$, our focus lies on the last two terms of (28), where $U_{b}$ and $U_{\eta}$ have the same impact on the optimal gap $G_{T}$. This implies that unbiased methods with large variance may ultimately achieve the same result as those biased ones at the same iteration step $T$. Hence in cases where variance dominates bias, methods with smaller variance are preferable.
On the other hand, when a sufficiently large iteration step $T$ is given and proper learning rate $\gamma$ is chosen, the upper bound of the optimal gap can be smaller than any given $\epsilon>0$ except for the parts of bias, which is examined in the following corollary.
Corollary 4.5.
Suppose that Assumptions 4.2 and 4.3 are satisfied and $\gamma=\min\left\{1/K,1/\mu,\epsilon\mu/(K\eta)\right\},\,T=\max\left\{(K/\mu)\ln(2G_{0}/\epsilon),K\eta/(\epsilon\mu^{2})\ln(2G_{0}{\epsilon})\right\}$. Then the optimal gap satisfies $G_{T}\leq\epsilon+U_{b}/(2\mu)$ for any $\epsilon>0$.
Proof.
It follows directly from (28) by scaling $\gamma$ and $T$. ∎
In this case, unbiased methods are always favorable, as they are able to converge to an arbitrarily small neighborhood of 0, while biased methods can only converge to a neighborhood of the $U_{b}$. Moreover, for biased method with $U_{b}\gg U_{\eta}$, the optimal gap for $t_{1}\gg t_{2}$ tends to be $G_{t_{1}}\approx G_{t_{2}}$, indicating that increasing the iteration step can be futile in attempting to reduce the optimal gap.
5 Numerical experiments
In this section, numerical experiments are conducted to compare the efficiency of two unbiased MLMC methods and a biased MLMC method for three models. To handle the curse of dimensionality, we propose to use low-dimensional summary statistics instead of the full data. However, this introduces additional bias, as detailed in [15]. In this paper, we do not look into the effects of using summary statistics for SNPE methods. We refer to Fearnhead and Prangle [15] for a semi-automatic method of constructing summary statistics in the context of ABC.
A toy example. Two-moon model was studied in Greenberg et al. [25]. For a given parameter $\theta\in\mathbb{R}^{2}$, the Two-moon simulator generates observations $x\in\mathbb{R}$ via
$a\sim U(-\frac{\pi}{2},\frac{\pi}{2})$ $,\quad r_{2}\sim\mathcal{N}(0.1,0.01^{2}),$
$p=(r_{2}\cos(a)+0.25,r_{2}\sin(a))$ $,\quad x=p+\left(-\frac{|\theta_{1}+\theta_{2}|}{\sqrt{2}},\frac{-\theta_{1}+\theta_{2}}{\sqrt{2}}\right).$
The intermediate variables $p$ follow a single crescent-shaped distribution, which is then shifted and rotated around the origin based on the parameter values of $\theta$. The absolute value $|\theta_{1}+\theta_{2}|$ contributes to the emergence of a second crescent in the posterior distribution. We choose a uniform prior over the square $[-1,1]^{2}$ to perform the inference.
Lotka-Volterra model. This model describes the continuous time evolution of a population of predators interacting with a population of prey using a stochastic Markov jump process. The model describe that, the birth of a predator at a rate $\exp(\theta_{1})XY$, resulting in an increase of $X$ by one; The death of a predator at a rate proportional to $\exp(\theta_{2})X$, leading to a decrease of $X$ by one; The birth of a prey at a rate proportional to $\exp(\theta_{3})Y$, resulting in an increase of $Y$ by one; The consumption of a prey by a predator at a rate proportional to $\exp(\theta_{4})XY$, leading to a decrease of $Y$ by one. Following the experimental details outlined in [49], we initialize the predator and prey populations as $X=50$ and $Y=100$, respectively. We conduct simulations of the Lotka-Volterra model using the Gillespie algorithm [21] over a duration of 30 time units. We recorded the populations at intervals of 0.2 time units, resulting in time series data sets, each consisting of 151 values. The resulting summary statistics $S(x)$ are represented as a 9-dimensional vector, which includes the following time series features: the logarithm of the mean of each time series, the logarithm of the variance of each time series, the auto-correlation coefficient of each time series at lags of 0.2 and 0.4 time units, and the cross-correlation coefficient between the two time series. In our experiments, the prior distribution of the parameters is set to $\mathcal{U}(-5,2)^{4}$, and we generate the ground truth posterior with SMC-ABC [4], which is more costly than methods in this paper.
M/G/1 queue model. The M/G/1 queue model [54] describes a single server’s processing of a queue of continuously arriving jobs. Define $I$ as the total number of jobs that needs to be processed, and denote by $s_{i}$ the processing time required for job $i$. Let $v_{i}$ be the job’s arrival time in the queue, and $d_{i}$ be the job’s departure time from the queue. They satisfy the following conditions
$s_{i}\sim\mathcal{U}(\theta_{1},\theta_{1}+\theta_{2}),\quad v_{i}-v_{i-1}\sim\mathrm{Exp}(\theta_{3}),\quad d_{i}-d_{i-1}=s_{i}+\max(0,v_{i}-d_{i-1}).$
In our experiments, we set $I=50$ jobs and the summary statistics $S(x)$ has been selected as the logarithm of 0th, 25th, 50th, 75th and 100th percentiles of the set of inter-departure times. The prior distribution of the parameters is
$\theta_{1}\sim\mathcal{U}(0,10),\quad\theta_{2}\sim\mathcal{U}(0,10),\quad\theta_{3}\sim\mathcal{U}(0,1/3),$
[…]
Our numerical experiments are performed on a computer equipped with a single GeForce RTX 2080s GPU and an i9-9900K CPU. The training and inference processes of the model are primarily implemented using the Pytorch package in Python.
In the training process, we simulate $N=1000$ samples in each round, and with $R=20$ rounds in total. In each round, we randomly pick 5% of the newly generated samples $\theta$ and their corresponding $x$ values as validation data. We follow the early stop criterion proposed by [49], which terminates the training if the loss value on the validation data does not decrease after 20 epochs in a single round. For the optimizer, we use Adam [36] with a batch size of 100, a learning rate of $1\times 10^{-4}$, and a weight decay of $1\times 10^{-4}$.
In this paper, we employ neural spline flows (NSFs) [14] as the conditional density estimator, which consists of 8 layers. Each layer is constructed using two residual blocks with 50 units and ReLU activation function. With 10 bins in each monotonic piecewise rational-quadratic transform, and the tail bound is set to 20.
We compare the performance of RU-MLMC, GRR-MLMC, and TGRR-MLMC on the Two-moon, Lotka-Volterra, and M/G/1 models. To assess the similarity between the approximate posterior distribution $q_{F(x_{o},\phi)}(\theta)$ and the true posterior distribution $p(\theta|x_{o})$ given the observed data, we employ maximum mean discrepancy (MMD) [25, 26, 31, 49] and classifier two-sample tests (C2ST) [12, 28, 40] as discriminant criteria. Additionally, we use log median distance (LMD) to measure the distance between $x_{o}$ and $x$ drawn from $p(x|\theta)$, where $\theta$ is sampled from $q_{F(x_{o},\phi)}(\theta)$. In cases where the true posterior distribution $p(\theta|x_{o})$ is intractable even with knowledge of the sample generation process, we resort to the negative log probability (NLOG) [14, 25, 31, 48, 49] of the true parameters $\theta^{*}$; in this scenario, the observation $x_{o}$ is sampled from $p(x|\theta^{*})$. It is important to note that lower values for all the mentioned indicators are favorable.
For the setting of MLMC methods, we have chosen $M_{0}=8$ and $\overline{m}=4$ for TGRR-MLMC. The base level for GRR and TGRR has been set to ${\underline{m}}=2$.
We now discuss the choice of the hyperparameter $\alpha$ for the geometric distribution of the level $L$. Unlike previous work where $\alpha$ is chosen to minimize the average cost $\mathrm{Cost}^{\mathrm{RU}}$ [22, 23, 30], based on experiment results in Figure 2, in addition to average cost, variance is another crucial factor that demands our attention. These two indicators collectively hold great significance in our evaluation and optimization processes. In order to reduce the variance of this estimate, one direct approach is to take average of different iterations. The asymptotic inefficiency [35] defined as $H_{\mathrm{RU}}:=\mathrm{Var}[V^{\mathrm{RU}}]\times\mathrm{Cost}_{\mathrm{RU}}$ remains a constant whether iteration average is taken. It would be effective to decrease variance by raising cost if $\alpha$ is chosen to minimize this quantity. Utilizing Theorem 3.5, we find an upper bound for the asymptotic inefficiency
$H_{\mathrm{RU}}\lesssim\frac{M_{0}2^{\alpha+r_{2}}}{\left(2^{\alpha}-2\right)\left(2^{r_{2}}-2^{\alpha}\right)}.$ (29)
As a result, the optimal $\alpha$ minimizing the upper bound is $\alpha_{\mathrm{RU}}^{*}=(r_{2}+1)/2\in(1,r_{2})$. When ${\underline{m}}=0$ in GRR-MLMC, which degenerates to RR-MLMC, the constants given in Theorem 3.6 do not involve the leverage of Theorem 3.4. Therefore, these constants are the same as those in Theorem 3.5, which yields an uniform upper bound for $H_{\mathrm{RR}}$. Therefore, we reach the same conclusion for $\alpha^{\mathrm{RR}}_{\mathrm{opt}}$. From this point of view, we can see that RR-MLMC actually trades lower variance for higher average cost compared with RU-MLMC. We apply the same procedure for other MLMC methods, the corresponding optimal $\alpha$ are presented in Table 1.
Table 1: The value of alpha minimizing the upper bound of asymptotic inefficiency
Method RU-MLMC GRR-MLMC TGRR-MLMC
$({\underline{m}}=2)$ $({\overline{m}}=4)$
$\alpha\ (1<\alpha<r_{2})$ 1.4 1.209 1.673
Since the proposal distributions of each round are distinct, the hyperparameter $r_{2}$ in Theorems 3.3 and 3.2 differs in each round. To determine the value of $r_{2}$ for each round, except for the first two, one may take the value from the previous round. To address this issue, we have conducted 50 training processes and sequentially performed linear regression, a universal value is then selected as $r_{2}=1.8$.
A
B
C
Figure 3: Density plot for RU-MLMC, GRR-MLMC and TGRR-MLMC A. Two-moon, from left to right: available ground truth, RU-MLMC, GRR-MLMC and TGRR-MLMC. B. Lotka-Volterra, from left to right: ground truth simulated with SMC-ABC [4], RU-MLMC, GRR-MLMC and TGRR-MLMC. C. M/G/1 queue model, the setting is the same with Lotka-Volterra.
A
B
C
Figure 4: Performance of RU-MLMC, GRR-MLMC and TGRR-MLMC A. Two-moon, B. Lotka-Volterra C. M/G/1 queue model, blue, green, and red correspond to RU-MLMC, GRR-MLMC, and TGRR-MLMC respectively.
We present our results in Figures 3 and 4. It is observed that both unbiased methods are inferior to the biased one in some cases. When comparing RU-MLMC with GRR-MLMC in the case of the Lotka-Volterra and M/G/1 queue models, we find that the unbiased method can greatly benefit from variance reduction. In the case of Lotka-Volterra, when comparing TGRR and GRR methods, we conclude that the gradient information $\Delta\rho_{\phi,\ell}$ where $\ell\geq\overline{m}$ not only does not contribute to the overall performance, but also make it worse when measured with C2ST. As suggested by Theorem 4.4, this could be due to the domination of variance, where the effect of reducing variance is superior to reducing bias.
In conclusion, in the case of SNPE where the complex model of $q_{F(x,\phi)}(\theta)$ is used for density estimation, variance tends to dominate the bias as it is suggested by Theorem 4.4, indicating that excessive variance could seriously affects the training process of the density estimator. Instead of seeking for unbiasedness, one should try to strike a balance between bias, average computational cost, and variance of the gradient in this case.
It is worth noting that in the case of the Two-moon problem, RU-MLMC method deliver comparable performance to the other two methods with lower computational cost and time requirements. Rather than employing any of these methods indiscriminately, it is crucial to undertake a comprehensive and thoughtful analysis of the specific problem at hand.
6 Concluding remarks
We develop a series of efficient nested MLMC methods, which are improvement over the unbiased MLMC methods. The latter suffer from excessive variance at the cost of unbiasedness, which, based on the convergence analysis, severely affects the method’s performance. Through experiments on standard benchmark for likelihood-free estimation and theoretical analysis, we valid the inferior performance of unbiased methods when variance dominates bias. In this paper, our choice for hyper-parameter $\alpha$ deviates from the mainstream choice. While values of $\alpha$ closer to $r_{2}$ is favored, aiming at cost minimization, given the fact that variance is the major focus, our focus lies on the asymptotic inefficiency by considering both variance and average cost. However, given that a thorough validation for this strategy through ablation experiments is still lacking, its improvement is still unclear. Quasi-Monte Carlo (QMC) and Randomized Quasi-Monte Carlo (RQMC) are well known variance reduction techniques, its application in nested MLMC methods has already been investigated in [30]. Given the high sensitivity of this problem to variance, using RQMC is likely to be beneficial, and existing results [30] can be appropriately utilized.
Acknowledgments
The work was supported by the National Natural Science Foundation of China grant 12071154, the Guangdong Basic and Applied Basic Research Foundation grant 2021A1515010275.
Appendix A Supplementary proofs
A.1 Proof for Theorem 3.3
Theorem A.1.
Assume that there exists $s>2$ that
$\underset{x,\phi,\theta}{\sup}\|\nabla_{\phi}\log g_{\phi}(x,\theta)\|_{\infty}=M_{max}<\infty,\,\underset{\phi}{\sup}\ \mathbb{E}\left[\left|\frac{g_{\phi}(x,\theta)}{Z(x,\phi)}\right|^{s}\right]<\infty,$
then we have
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\right]$ $\leq\frac{16(2^{r_{2}}+1)(E_{1}+E_{2})H(2r)+9E_{3}H(s)1\{s<4\}}{(M_{0}2^{\ell})^{r_{2}}},$ (30)
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|\right]$ $\leq\frac{(6\sqrt{2}d^{1/4}M^{1/2}_{\max}+4d^{1/2}M_{\max})C_{2}H(2)}{M_{\ell}},$ (31)
where
$E_{1}$ $=2^{6-1/2r-r_{2}}dM^{2}_{\max}C_{2r},$
$E_{2}$ $=2^{4-2r}dM^{2}_{\max}C_{2r},$
$E_{3}$ $=2^{1+3s/2}dM^{2}_{\max}C_{s},$
$H(x)$ $=\underset{\phi\in\Phi}{\sup}\ \mathbb{E}\left[\left|\frac{g_{\phi}(x,\theta)}{Z(x,\phi)}\right|^{x}\right]+1<\infty,\quad x\leq s$
[…]
where the latter one requires the additional assumption that the supremum of the ratio $a/b$ is bounded. Next, given for any $u>2$, for $|x-1|<1/2$ we arrive that for any $\tau\in(0,4]$:
$(x-1)^{\tau}\leq 2^{4-\min(u,4)}\left|x-1\right|^{\min(u,4)-(4-\tau)}.$ (33)
This proof follows [23], since the implicit constant is of interest, we still present it here for the sake of completeness.
Proof.
Consider the following event $\mathcal{A}$
$\mathcal{A}:=\left\{\left|S^{(a)}_{\ell-1}\right|>\frac{1}{2}\right\}\bigcup\left\{\left|S^{(b)}_{\ell-1}\right|>\frac{1}{2}\right\},$
where $S^{(a)}_{\ell-1}:=g^{(a)}_{\phi,M_{\ell-1}}(x)/Z(x,\phi)-1$, which is the sample mean of $M_{\ell-1}$ variable $\mathfrak{s}=g_{\phi}(x,\theta)/Z(x,\phi)-1$ with zero mean. Similarly, we define $S^{(b)}_{\ell-1}$ and $S_{\ell}:=g_{\phi,M_{\ell}}(x)/Z(x,\phi)-1$ and arrive at the following decomposition
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\right]=\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\mathbf{1}_{\mathcal{A}}\right]+\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\mathbf{1}_{\mathcal{A}^{c}}\right].$ (34)
We then need to develop a proper upper bound for this two terms. Denote the dimension of the $\phi$ as $d$, which is always finite, we leverage (32) to obtain
$\left\|\rho_{\phi,M_{\ell}}\right\|^{2}_{2}\leq\left\|\frac{\nabla_{\phi}g_{\phi}(x,\theta)}{g_{\phi}(x,\theta)}\right\|^{2}+\left\|\frac{\sum_{j=1}^{M_{\ell}}\nabla_{\phi}g_{\phi}(x,\theta_{j}^{\prime})}{\sum_{j=1}^{M_{\ell}}g_{\phi}(x,\theta_{j}^{\prime})}\right\|^{2}_{2}\leq 2dM_{\max}^{2},$
giving us a rough bound of the $\ell_{2}$ norm of $\rho_{\phi,M_{\ell}}$. That is
$\max\left\{\left\|\rho^{(a)}_{\phi,M_{\ell-1}}\right\|_{2}^{2},\left\|\rho^{(b)}_{\phi,M_{\ell-1}}\right\|_{2}^{2}\right\}\leq 2dM^{2}_{\max}.$
Then for $\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}$, we apply Jensen’s inequality to develop an upper bound
$\left\|\Delta\rho_{\phi,\ell}\right\|^{2}_{2}$ $\leq\left(\left\|\rho_{\phi,M_{\ell}}\right\|_{2}+\frac{\left\|\rho^{(a)}_{\phi,M_{\ell-1}}\right\|_{2}}{2}+\frac{\left\|\rho^{(b)}_{\phi,M_{\ell-1}}\right\|_{2}}{2}\right)^{2}$
$\leq 3\left\|\rho_{\phi,M_{\ell}}\right\|_{2}^{2}+\frac{3}{4}\left\|\rho^{(a)}_{\phi,M_{\ell-1}}\right\|_{2}^{2}+\frac{3}{4}\left\|\rho^{(b)}_{\phi,M_{\ell-1}}\right\|_{2}^{2}\leq 9dM^{2}_{\max}.$
Then for the first term in (34) we have
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\mathbf{1}_{\mathcal{A}}\right]\leq\|\Delta\rho_{\phi,\ell}\|_{2}^{2}\mathbb{P}\left[\mathcal{A}\right]\leq 9dM^{2}_{\max}\mathbb{P}\left[\mathcal{A}\right].$ (35)
As for $\mathbb{P}[\mathcal{A}]$, applying Lemma 3.1 to have
$\mathbb{P}\left[\mathcal{A}\right]\leq\mathbb{P}\left[\left|S^{(a)}_{\ell-1}\right|>\frac{1}{2}\right]+\mathbb{P}\left[\left|S^{(b)}_{\ell-1}\right|>\frac{1}{2}\right]\leq\frac{2^{3s/2+1}C_{s}}{(M_{0}2^{\ell})^{s/2}}H(s),$
and it is proved that $\mathbb{E}[\|\Delta\rho_{\phi,\ell}\|_{2}^{2}\mathbf{1}_{\mathcal{A}}]$ is of order $2^{-(s/2)\ell}$. For the second term $\mathbb{E}[\|\Delta\rho_{\phi,\ell}\|_{2}^{2}\mathbf{1}_{\mathcal{A}^{c}}]$ in (34), utilizing the antithetic property to attain the following identity
$\Delta\rho_{\phi,\ell}$ $=\frac{1}{2}\left(\nabla g^{(a)}_{\phi,M_{\ell-1}}(x)-\nabla_{\phi}Z(x,\phi)\right)\left(\frac{1}{g^{(a)}_{\phi,M_{\ell-1}}(x)}-\frac{1}{Z(x,\phi)}\right)$
$+\frac{1}{2}\left(\nabla g^{(b)}_{\phi,M_{\ell-1}}(x)-\nabla_{\phi}Z(x,\phi)\right)\left(\frac{1}{g^{(b)}_{\phi,M_{\ell-1}}(x)}-\frac{1}{Z(x,\phi)}\right)$
$-\left(\nabla g_{\phi,M_{\ell}}(x)-\nabla_{\phi}Z(x,\phi)\right)\left(\frac{1}{g_{\phi,M_{\ell}}(x)}-\frac{1}{Z(x,\phi)}\right)$
$+\frac{1}{2}\frac{\nabla_{\phi}Z(x,\phi)}{g^{(a)}_{\phi,M_{\ell-1}}(x)}\left(\frac{g^{(a)}_{\phi,M_{\ell-1}}(x)}{Z(x,\phi)}-1\right)^{2}+\frac{1}{2}\frac{\nabla_{\phi}Z(x,\phi)}{g^{(b)}_{\phi,M_{\ell-1}}(x)}\left(\frac{g^{(b)}_{\phi,M_{\ell-1}}(x)}{Z(x,\phi)}-1\right)^{2}$
$-\frac{\nabla_{\phi}Z(x,\phi)}{g_{\phi,M_{\ell}}(x)}\left(\frac{g_{\phi,M_{\ell}}(x)}{Z(x,\phi)}-1\right)^{2}.$ (36)
Now on event $\mathcal{A}^{c}$, we have $1/{g^{(a)}_{\phi,M_{\ell-1}}(x)}\leq 2/{Z(x,\phi)},1/{g^{(b)}_{\phi,M_{\ell-1}}(x)}\leq 2/{Z(x,\phi)}$. With simple algebra, we have $1/{g_{\phi,M_{\ell}}(x)}\leq 2/{Z(x,\phi)}$. Also we have $|S_{\ell}|<1/2$ on $\mathcal{A}^{c}$. Then an upper bound of the normalizing constant estimators on event $\mathcal{A}^{c}$ are given. We denote $F^{(a)}_{\ell-1}:=(\nabla g^{(a)}_{\phi,M_{\ell-1}}(x)-\nabla_{\phi}Z(x,\phi))/Z(x,\phi),\tilde{F}^{(a)}_{\ell-1}:=(\nabla g^{(a)}_{\phi,M_{\ell-1}}(x)-\nabla_{\phi}Z(x,\phi))/g^{(a)}_{\phi,M_{\ell-1}}$. Where the former one is also a mean of $M_{\ell-1}$ variable $\mathfrak{f}:=\nabla_{\phi}g_{\phi}(x,\theta)-\nabla_{\phi}Z(x,\phi)/Z(x,\phi)$ with zero mean. Similarly, we define $F^{(b)}_{\ell-1},\tilde{F}^{(b)}_{\ell-1}$ and $F_{\ell}:=(\nabla g_{\phi,M_{\ell}}(x)-\nabla_{\phi}Z(x,\phi))/Z(x,\phi),\tilde{F}_{\ell}:=(\nabla g_{\phi,M_{\ell}}(x)-\nabla_{\phi}Z(x,\phi))/g_{\phi,M_{\ell}}(x)$. Directly applying Jensen’s inequality on (36) to have
$\left\|\Delta\rho_{\phi,\ell}\right\|^{2}_{2}$ $\leq 2\|\tilde{F}^{(a)}_{\ell-1}\|_{2}^{2}(S^{(a)}_{\ell-1})^{2}+2\|\tilde{F}^{(b)}_{\ell-1}\|_{2}^{2}(S^{(b)}_{\ell-1})^{2}+4\|\tilde{F}_{\ell}\|_{2}^{2}(S_{\ell})^{2}$
$+2\left\|\frac{\nabla_{\phi}Z(x,\phi)}{g^{(a)}_{\phi,M_{\ell-1}}(x)}\right\|_{2}^{2}(S^{(a)}_{\ell-1})^{4}+2\left\|\frac{\nabla_{\phi}Z(x,\phi)}{g^{(b)}_{\phi,M_{\ell-1}}(x)}\right\|_{2}^{2}(S^{(b)}_{\ell-1})^{4}$
$+4\left\|\frac{\nabla_{\phi}Z(x,\phi)}{g_{\phi,M_{\ell}}(x)}\right\|_{2}^{2}(S_{\ell})^{4}.$ (37)
With the upper bound derived before, we arrive at the following:
$\left\|\Delta\rho_{\phi,\ell}\right\|^{2}_{2}$ $\leq 8\|F^{(a)}_{\ell-1}\|_{2}^{2}(S^{(a)}_{\ell-1})^{2}+8\|F^{(b)}_{\ell-1}\|_{2}^{2}(S^{(b)}_{\ell-1})^{2}+16\|F_{\ell}\|_{2}^{2}(S_{\ell})^{2}$
$+8\left\|\frac{\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|_{2}^{2}\left((S^{(a)}_{\ell-1})^{4}+(S^{(b)}_{\ell-1})^{4}+2(S_{\ell})^{4}\right),$ (38)
where each term is some power of a mean of $\mathfrak{f}$ or $\mathfrak{s}$ enabling one to use Lemma 3.1 to derive their upper bound. To obtain the order of $\mathbb{E}[\|\Delta\rho_{\phi,l}\|_{2}^{2}\mathbf{1}_{\mathcal{A}^{c}}]$, we begin by applying (32) to yield
$\left\|\frac{\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|_{2}^{2}=\left\|\frac{\mathbb{E}[\nabla_{\phi}g_{\phi}(x,\theta)]}{\mathbb{E}[g_{\phi}(x,\theta)]}\right\|_{2}^{2}\leq\underset{x,\theta,\phi}{\mathrm{sup}}\left\|\frac{\nabla_{\phi}g(x,\theta)}{g_{\phi}(x,\theta)}\right\|_{2}^{2},$ (39)
then it would be sufficient to derive an upper bound for the expectation of the $\|F_{\ell}\|_{2}^{2}(S_{\ell})^{2}$ and $(S_{\ell})^{4}$ in (38) on event $\mathcal{A}^{c}$ due to the symmetry. We firstly start from $\|F_{\ell}\|_{2}^{2}(S_{\ell})^{2}$, applying (33) to obtain
$\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2}\left(S_{\ell}\right)^{2}\textbf{1}_{\mathcal{A}^{c}}\right]$ $\leq\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2}\times 2^{4-2r}\left|S_{\ell}\right|^{{2r-2}}\right]$
$\leq\left(\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2r}\right]\right)^{1/r_{2}}2^{4-2r}\left(\mathbb{E}\left[\left|S_{\ell}\right|^{{2r}}\right]\right)^{1-1/r_{2}}.$ (40)
To obtain the desired upper bound, we handle with these two moments separately. For $\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2r}\right]$, with Jensen’s inequality and Lemma 3.1, we have
$\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2r}\right]\leq\frac{C_{2r}}{M_{\ell}^{r_{2}}}\mathbb{E}\left[\|\mathfrak{f}\|^{2r}_{2}\right],$
then for $\mathbb{E}\left[\|\mathfrak{f}\|^{2r}_{2}\right]$ on $\mathcal{A}^{c}$, applying Jensen’s inequality
$\mathbb{E}\left[\|\mathfrak{f}\|^{2r}_{2}\right]$ $=\mathbb{E}\left[\left\|\frac{\nabla_{\phi}g_{\phi}(x,\theta)-\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|^{2r}_{2}\right]$
$\leq d^{r_{2}-1}2^{2r-1}\mathbb{E}\left[\left\|\frac{\nabla_{\phi}g_{\phi}(x,\theta)}{g_{\phi}(x,\theta)}\cdot\frac{g_{\phi}(x,\theta)}{Z(x,\phi)}\right\|_{2r}^{2r}+\left\|\frac{\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|_{2r}^{2r}\right].$
Then an upper bound for $\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2r}\right]$ in (40) is presented as
$\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2r}\right]\leq\frac{2^{2r-1}d^{r_{2}}M^{2r}_{\max}C_{2r}}{M_{\ell}^{r_{2}}}H(2r).$
For the upper bound of $\mathbb{E}\left[\left|S_{\ell}\right|^{{2r}}\right]$ in (40), we directly apply Lemma 3.1 to have
$\mathbb{E}\left[\left|S_{\ell}\right|^{{2r}}\right]\leq\frac{C_{2r}}{M_{\ell}^{r_{2}}}H(2r),$
directly plugging them into (40) to obtain the desired upper bound
$\mathrm{RHS}=\left(2^{2r-1}d^{r_{2}}M^{2r}_{\max}C_{2r}\right)^{1/r_{2}}2^{4-2r}C^{1-1/{r_{2}}}_{2r}\frac{H(2r)}{(M_{0}2^{\ell})^{r_{2}}}=\frac{E_{1}H_{2r}}{(M_{0}2^{\ell})^{r_{2}}}.$
With similar treatment for $\mathbb{E}[\|F^{(a)}_{\ell-1}\|_{2}^{2}(S^{(a)}_{\ell-1})^{2}1_{\mathcal{A}^{c}}]$ and $\mathbb{E}[\|F^{(b)}_{\ell-1}\|_{2}^{2}(S^{(b)}_{\ell-1})^{2}1_{\mathcal{A}^{c}}]$, we then have a uniform upper bound as $E_{1}H_{2r}2^{r_{2}}/(M_{0}2^{\ell})^{r_{2}}$.
As for the $\mathbb{E}[(S_{\ell})^{4}1_{\mathcal{A}^{c}}]$, the sixth term in (38), note that
$\mathbb{E}[(S_{\ell})^{4}1_{\mathcal{A}^{c}}]\leq 2^{4-2r}\mathbb{E}\left[\left|S_{\ell}\right|^{2r}\right]\leq 2^{4-2r}\frac{C_{2r}}{M_{\ell}^{r_{2}}}H(2r)$
Combining these upper bounds, we arrive at:
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\right]\leq\frac{16(2^{r_{2}}+1)(E_{1}+E_{2})H(2r)+9E_{3}H(s)1\{s<4\}}{(M_{0}2^{\ell})^{r_{2}}}.$
For the second required upper bound, we follow from (36) that
$\left\|\Delta\rho_{\phi,\ell}\right\|_{2}$ $\leq\|F^{(a)}_{\ell-1}\|_{2}|S^{(a)}_{\ell-1}|+\|F^{(b)}_{\ell-1}\|_{2}|S^{(b)}_{\ell-1}|+2\|F_{\ell}\|_{2}|S_{\ell}|$ (41)
$+\left\|\frac{\nabla Z_{\phi}(x,\phi)}{Z(x,\phi)}\right\|_{2}\left((S^{(a)}_{\ell-1})^{2}+(S^{b}_{\ell})^{2}+2(S_{\ell})^{2}\right),$ (42)
the same decomposition on event $A$ is also studied. We apply Lemma 3.1 and (39) to derive an upper bound for the second row. For the first row, it would be sufficient to find an upper bound for $\mathbb{E}[\|F_{\ell}\|_{2}|S_{\ell}|1_{\mathcal{A}^{c}}]$, applying Holder’s inequality to yield that
$\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}\left|S_{\ell}\right|1_{A^{c}}\right]\leq\left(\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2}\right]\right)^{1/2}\left(\mathbb{E}\left[\left|S_{\ell}\right|^{2}\right]\right)^{1/2}.$
We then utilize Lemma 3.1 to have
$\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2}1_{\mathcal{A}^{c}}\right]\leq\frac{2d^{1/2}M_{\max}C_{2}}{M_{\ell}}H(2),\quad\mathbb{E}[\left|S_{\ell}\right|^{2}1_{\mathcal{A}^{c}}]\leq\frac{C_{2}}{M_{\ell}}H(2).$
Therefore the upper bound for $\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}\left|S_{\ell}\right|1_{A^{c}}\right]$ is $\sqrt{2}d^{1/4}M^{1/2}_{\max}C_{2}H(2)/M_{\ell}$. Finally we have the following upper bound for $\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|\right]$
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|\right]\leq\frac{(6\sqrt{2}d^{1/4}M^{1/2}_{\max}+4d^{1/2}M_{\max})C_{2}H(2)}{M_{\ell}}.$
A.2 Proof for Theorem 3.4
Theorem A.2.
Under the assumptions of Theorem 3.3, we have
$\mathrm{Var}\left[\rho_{\phi,M_{\ell}}\right]$ $\leq\frac{(16dM_{\max}^{2}+32d^{1/2}M_{\max})C_{2}H(2)}{M_{0}2^{\ell}}$
$+\frac{\left(8dM^{2}_{\max}+24\sqrt{2}d^{3/4}M^{3/2}_{\max}+36d^{1/2}M_{\max}\right)C^{2}_{2}H^{2}_{2}}{(M_{0}2^{\ell})^{2}}.$ (43)
Proof.
By Jensen’s equality, we have
$\mathrm{Var}\left[\rho_{\phi,M_{\ell}}\right]$ $=\mathbb{E}\left[\left\|\rho_{\phi,M_{\ell}}-\mathbb{E}\left[\rho_{\phi,M_{\ell}}\right]\right\|_{2}^{2}\right]$
$=\mathbb{E}\left[\left\|\rho_{\phi,M_{\ell}}-\rho_{\phi}+\rho_{\phi}-\mathbb{E}\left[\rho_{\phi,M_{\ell}}\right]\right\|_{2}^{2}\right]$
$\leq 2\mathbb{E}\left[\left\|\rho_{\phi,M_{\ell}}-\rho_{\phi}\right\|_{2}^{2}\right]+2\mathbb{E}\left[\left\|\rho_{\phi}-\mathbb{E}\left[\rho_{\phi,M_{\ell}}\right]\right\|_{2}^{2}\right].$ (44)
For the first term of (43), we begin with
$\rho_{\phi,M_{\ell}}-\rho_{\phi}=$ $\left(-\nabla_{\phi}\log g_{\phi}(x,\theta)+\nabla_{\phi}\log\hat{Z}_{M_{\ell}}(x,\phi)\right)$
$-\left(-\nabla_{\phi}\log g_{\phi}(x,\theta)+\nabla_{\phi}\log Z(x,\phi)\right)$
$=$ $\ \frac{\nabla g_{\phi,M_{\ell}}(x)}{g_{\phi,M_{\ell}}(x)}-\frac{\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)},$
and similarly, we consider the event
:=$\mathcal{B}:=\left\{\left|\frac{g_{\phi,M_{\ell}}(x)}{Z(x,\phi)}-1\right|>\frac{1}{2}\right\},$
and the similar decomposition used for event $\mathcal{A}$ in the proof of Theorem 3.3 is studied. Firstly we derive a rough bound for $\left\|\rho_{\phi,M_{\ell}}-\rho_{\phi}\right\|_{2}^{2}$ with (32)
$\left\|\rho_{\phi,M_{\ell}}-\rho_{\phi}\right\|_{2}^{2}\leq 2\left\|\frac{\nabla g_{\phi,M_{\ell}}(x)}{g_{\phi,M_{\ell}}(x)}\right\|_{2}^{2}+2\left\|\frac{\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|_{2}^{2}\leq 4dM_{\max}^{2}.$
Then for event $\mathcal{B}$, by Lemma 3.1 again, we obtain
$\mathbb{E}\left[\left\|\rho_{\phi,M_{\ell}}-\rho_{\phi}\right\|_{2}^{2}\mathbf{1}_{B}\right]\leq 4dM_{\max}\mathbb{P}[B]\leq\frac{2^{s+2}dM_{\max}C_{s}}{(M_{0}2^{\ell})^{s/2}}H(s).$
On event $\mathcal{B}^{c}$, we have
$\left\|\rho_{\phi,M_{\ell}}-\rho_{\phi}\right\|^{2}_{2}$ $=\left\|\frac{\nabla g_{\phi,M_{\ell}}(x)-\nabla_{\phi}Z(x,\phi)}{g_{\phi,M_{\ell}}(x)}+\frac{\nabla_{\phi}Z(x,\phi)}{g_{\phi,M_{\ell}}(x)}\left(\frac{g_{\phi,M_{\ell}}(x)}{Z(x,\phi)}-1\right)\right\|^{2}_{2}$
$\leq 8\left\|F_{\ell}\right\|_{2}^{2}+8\left\|\frac{\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|^{2}_{2}S_{\ell}^{2},$
and
$\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2}1_{\mathcal{B}^{c}}\right]\leq\frac{2d^{1/2}M_{\max}C_{2}}{M_{\ell}}H(2),\quad\mathbb{E}[\left|S_{\ell}\right|^{2}1_{\mathcal{B}^{c}}]\leq\frac{C_{2}}{M_{\ell}}H(2).$
These result in the following upper bound
$\mathbb{E}\left[\left\|\rho_{\phi,M_{\ell}}-\rho_{\phi}\right\|_{2}^{2}\right]\leq\frac{(16d^{1/2}M_{\max}+8dM_{\max}^{2})C_{2}H(2)}{M_{0}2^{\ell}}.$
For the second term of (43), we have
$\left\|\mathbb{E}\left[\rho_{\phi,M_{\ell}}\right]-\rho_{\phi}\right\|_{2}=\left\|\sum_{\ell^{\prime}=\ell+1}^{\infty}\mathbb{E}\left[\Delta\rho_{\phi,\ell^{\prime}}\right]\right\|_{2}\leq\sum_{\ell^{\prime}=\ell+1}^{\infty}\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell^{\prime}}\right\|_{2}\right],$
then applying Theorem 3.3 to yield the desired result
$\left\|\mathbb{E}\left[\rho_{\phi,M_{\ell}}\right]-\rho_{\phi}\right\|_{2}\leq\frac{(3\sqrt{2}d^{1/4}M^{1/2}_{\max}+2d^{1/2}M_{\max})C_{2}H(2)}{M_{0}2^{\ell}}.$ (45)
A.3 Proof of Theorem 3.6
Proof.
Applying Theorem 3.3, we have the following upper bound for the variance of the gradient estimators of GRR
$\mathrm{Var}\left[V_{\mathrm{GRR}}\right]$ $=\mathrm{Var}\left[\rho_{\phi,M_{{\underline{m}}}}\right]+\sum_{\ell={\underline{m}}+1}^{\infty}w_{\ell}\sum_{\ell^{\prime}={\underline{m}}+1}^{\ell}(1-p)^{-2\ell^{\prime}}\mathrm{Var}\left[\Delta\rho_{\phi,\ell^{\prime}}\right]$
$\leq\frac{B_{1}}{2^{\underline{m}}}+\frac{B_{2}}{2^{2\underline{m}}}+\sum_{\ell={\underline{m}}+1}^{\infty}w_{\ell}\sum_{\ell^{\prime}={\underline{m}}+1}^{\ell}(1-p)^{-2\ell^{\prime}}B_{3}2^{-r_{2}\ell^{\prime}}$
$=\frac{B_{1}}{2^{\underline{m}}}+\frac{B_{2}}{2^{2\underline{m}}}+B_{3}\left(\sum_{\ell={\underline{m}}+1}^{\infty}w_{\ell}\frac{2^{(2\alpha-r_{2})({\underline{m}}+1)}(1-2^{(2\alpha-r_{2})(\ell-{\underline{m}})})}{1-2^{2\alpha-r_{2}}}\right)$
$=\frac{B_{1}}{2^{\underline{m}}}+\frac{B_{2}}{2^{2\underline{m}}}+B_{3}\left(\sum_{\ell={\underline{m}}+1}^{\infty}(1-p)^{\ell}p\frac{2^{(2\alpha-r_{2})({\underline{m}}+1)}-2^{(2\alpha-r_{2})(\ell+1)}}{1-2^{2\alpha-r_{2}}}\right)$
$=\frac{B_{1}}{2^{\underline{m}}}+\frac{B_{2}}{2^{2\underline{m}}}+B_{3}\left(\frac{1-2^{-\alpha}}{1-2^{2\alpha-r_{2}}}\left(\frac{2^{(\alpha-r_{2})({\underline{m}}+1)}}{1-2^{-\alpha}}-\frac{2^{2\alpha-r_{2}}2^{(\alpha-r_{2})({\underline{m}}+1)}}{1-2^{\alpha-r_{2}}}\right)\right)$
$\leq\frac{B_{1}+B_{2}}{2^{\underline{m}}}+B_{3}\left(\frac{2^{(\alpha-r_{2})({\underline{m}}+1)}}{1-2^{\alpha-r_{2}}}\right)$
$\leq\frac{B}{2^{\underline{m}}}+B\left(\frac{2^{(\alpha-r_{2})({\underline{m}}+1)}}{1-2^{\alpha-r_{2}}}\right),$
where $B=\max\{B_{1}+B_{2},B_{3}\}$ with $B_{1},\ B_{2}$ and $B_{3}$ representing the constants in the upper bounds in Theorems 3.4 and 3.3. The proof for the upper bound of $\mathrm{Var}[U_{\mathrm{GRR}}]$ follows a similar procedure. The average cost is
$\mathrm{Cost}_{\mathrm{GRR}}$ $\propto M_{0}2^{{\underline{m}}}+\sum_{\ell={\underline{m}}+1}^{\infty}w_{\ell}\sum_{\ell^{\prime}={\underline{m}}+1}^{\ell}M_{0}2^{\ell^{\prime}}$
$=M_{0}2^{{\underline{m}}}+M_{0}(1-2^{-\alpha})\sum_{\ell={\underline{m}}+1}^{\infty}2^{-\alpha\ell}\left(2^{\ell+1}-2^{{\underline{m}}+1}\right)$
$=M_{0}2^{{\underline{m}}}+M_{0}(1-2^{-\alpha})\left(\frac{2\times 2^{(1-\alpha)({\underline{m}}+1)}}{1-2^{(1-\alpha)}}-\frac{2^{(1-\alpha)({\underline{m}}+1)}}{1-2^{-\alpha}}\right)$
$=M_{0}2^{\underline{m}}+M_{0}\frac{2^{(1-\alpha)({\underline{m}}+1)}}{1-2^{(1-\alpha)}},$
completing the proof. ∎